OpenStudy (anonymous):

How do I simplify rational exponents? Such as, (16^-3/5)^5/4

OpenStudy (jdoe0001):

keep in mind that \(\bf \large a^{\frac{{\color{blue} n}}{{\color{red} m}}} = \sqrt[{\color{red} m}]{a^{\color{blue} n}}\qquad \qquad a^{-{\color{red} n}} = \cfrac{1}{a^{\color{red} n}}\\ \quad \\ \quad \\ \left(16^{-\frac{3}{5}}\right)^{\frac{5}{4}}\implies 16^{-\frac{3}{5}\cdot \frac{5}{4}}\)

OpenStudy (jdoe0001):

that is \(\Large \bf \left(16^{-\frac{3}{5}}\right)^{\frac{5}{4}}\implies 16^{-\frac{3}{5}\cdot \frac{5}{4}}\)

OpenStudy (anonymous):

@jdoe0001 Thank you so much, I appreciate your help.

OpenStudy (jdoe0001):

yw